### v1

# Generating Fine-Grained Human Motions Using ChatGPT-Refined Descriptions

Abstract
Recently, significant progress has been made in text-based motion generation, enabling the generation of diverse and high-quality human motions that conform to textual descriptions. However, it remains challenging to generate fine-grained or stylized motions due to the lack of datasets annotated with detailed textual descriptions. By adopting a divide-and-conquer strategy, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for human motion generation. Specifically, we first parse previous vague textual annotation into fine-grained description of different body parts by leveraging a large language model (GPT-3.5). We then use these fine-grained descriptions to guide a transformer-based diffusion model. FG-MDM can generate fine-grained and stylized motions even outside of the distribution of the training data. Our experimental results demonstrate the superiority of FG-MDM over previous methods, especially the strong generalization capability. We will release our fine-grained textual annotations for HumanML3D and KIT.
Figure 1: FG-MDM can generate high-quality fine-grained and stylized motions by using fine-grained descriptions of different body parts.
1 Introduction
Human motion generation is an important research topic in communities of both computer vision and computer graphics. It aims to simulate and generate realistic human movements using computers. With the advancement of technologies such as virtual reality, augmented reality, and movie special effects, there is a growing demand for high-quality human motion generation. In recent years, quite a few innovative methods and techniques have emerged to tackle this challenging task [45]. Deep generative models, including GANs [20, 1], VAEs [9, 25, 24], and Diffusion models [35, 42, 16, 5], have been widely applied to human motion generation.
However, there is relatively less research on generating fine-grained motions, which usually mobilize the entire body and involve multiple actions, based on textual descriptions. In [3], Athanasiou et al. proposed a spatial composition framework named SINC for simultaneous action generation. It should be noted that SINC was designed for motion generation based on action labels, not using natural language. MotionDiffuse can achieve fine-grained controlling when generating motions using textual descriptions [42]. However, the controlling ability is largely limited by the scarcity of the training data. Likewise, there are very few research efforts devoted to stylized motion generation with natural language, which generates motions accompanied with rich styles and emotions. In [34], Tevet et al. proposed MotionCLIP to study stylized motion generation. However, only simple action classes with several style labels were considered. Stylized motions defined by textual descriptions were not investigated.
A challenging issue of current human motion generation is the scarcity of datasets. This is especially true for datasets that include fine-grained descriptions or stylized text annotations. Datasets, such as KIT [26], BABEL [27], and HumanML3D [9], provide rich motion information. Their textual descriptions instead are often simple and direct, lacking stylized and fine-grained descriptions. Therefore, compared to traditional human motion generation, generating fine-grained and stylized motions is more challenging due to the limited scale and diversity of existing motion capture datasets. In Figure 3 and Figure 4, we show some fine-grained and stylized motions generated by MDM [35] and MLD [5]. Although both approaches achieve impressive performance in human motion generation, they show poor generalization ability since these fine-grained and stylized motions are outside of the distribution of the training data. Recently, there have been attempts to leverage human mesh recovery approaches [36, 40, 39] to collect large-scale text pseudo-pose datasets [4]. As shown in Azadi et al. [4], the pre-training on such datasets can significantly improve the generalization to in-the-wild descriptions, but generating fine-grained and stylized motions remains challenging due to the lack of detailed descriptions.
Then, with only limited motion capture datasets available, can we still achieve fine-grained and stylized motion generation? For a textual description defining a fine-grained or stylized motion, is there any way to associate it with motions within the dataset? Note that the fine-grained and stylized motions might be outside of the distribution of the dataset. For a never-before-seen motion, the motion of the entire body is indeed unseen, but motions of specific body parts may be inside of the distribution of the dataset. Therefore, we can adopt a divide-and-conquer strategy. By re-annotating the motion for different body parts with fine-grained descriptions, we can associate these body parts with specific body parts within the dataset. For example, the stylized description “A person walks depressingly.” can be reformulated as “His arms hang heavily by his sides. His legs move slowly, taking short steps with little energy…”. Leg movement in this stylized motion may appear in “A person walks aimlessly and slowly.”. The arm movement may appear in “A person stands in a relaxed manner.”. In Figure LABEL:figshow, we give two examples. On one hand, adopting fine-grained descriptions allows the model to get how the fine-grained or stylized motion is performed in detail. On the other hand, re-annotating the motion for different body parts enables the model to handle motions beyond the scope of the dataset.
Although manually annotating fine-grained textual descriptions for body parts provides more accurate data, it requires a significant amount of manual work, resulting in huge cost. Fortunately, with the rapid development of large language models, OpenAI’s GPT series models [23], known for their outstanding natural language processing capabilities, have gained widespread attention worldwide. In [7], Gilardi et al. demonstrated that ChatGPT performs as well as human annotators in some tasks. In [15], Action-GPT explores the excellent capability of ChatGPT in expanding human action descriptions. However, the generated content tends to be excessively redundant. For our task, we carefully design a prompt that allows ChatGPT-3.5 to provide detailed but non-redundant transcriptions of text descriptions about the human motion. We then use this prompt and ChatGPT-3.5 to transcribe 44,970 short text descriptions from HumanML3D [9] and 6,353 text descriptions from KIT [26] for fine-grained and stylized motion generation. We will make these fine-grained transcriptions publicly available.
With these fine-grained descriptions, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for human motion generation. Specifically, we replace the original simple and direct text with ChatGPT-generated fine-grained descriptions of individual body parts to guide a transformer-based diffusion model. Following MDM [35], we encode the entire fine-grained description with CLIP [29] as a global token of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens. By adopting these two tokens, the model can pay attention to both the global and detailed information of human motions, thereby improving the accuracy and completeness of the denoising results.
Figure 2: The overall pipeline of FG-MDM. The model learns the denoising process of the diffusion model from the motion $x_{t}^{1:n}$ at time step $t$ to the clean motion $\hat{x}_{0}^{1:n}$, given the text condition. The input text is first paraphrased by ChatGPT into fine-grained descriptions $D_{1:k}$ for different parts of the body, where $k$ denotes the number of body parts. These descriptions are then fed into a pre-trained CLIP text encoder and projected, along with the time step $t$, onto input tokens $PT_{1:k}$ of the transformer. The overall fine-grained text is further encoded into a global input token $GL$, providing holistic information. In the sampling process of the diffusion model, an initial random noise $x_{T}^{1:n}$ is sampled, and then $T$ iterations are performed to generate the clean motion $\hat{x}_{0}^{1:n}$. At each sampling step $t$, guided by $PT_{1:k}$ and $GL$, the transfomer encoder predicts the clean motion $\hat{x}_{0}^{1:n}$ which is then noised back to $x_{t-1}^{1:n}$.
Our contributions are summarized as follows:
  • We present a novel framework that utilizes fine-grained descriptions of different body parts to guide the training of the human motion diffusion model. This framework enables the generation of fine-grained and stylized motions even in the absence of datasets annotated with fine-grained and stylized descriptions.
  • We carefully design a prompt that enables ChatGPT to convert short texts into detailed but non-redundant descriptions of different body parts. We then use this prompt to transcribe 44,970 texts from HumanML3D and 6,353 texts from KIT into fine-grained descriptions. We will make these fine-grained transcriptions publicly available.
  • We conduct a series of experiments to evaluate our model’s ability in not only fitting the training data but also generating fine-grained and stylized motions, i.e., the generalization capability.
2 Related Work
2.1 Human Motion Generation
Recent years have seen a great interest in human motion generation. Previous work has explored unconditional generative models [44, 28] as well as generative models using various input conditions, such as text [9, 35, 5], prior motion [22, 10], action class [8, 24], and music [18, 37, 17]. In this paper, we focus on text-to-motion generation. Early work usually addressed the text-to-motion task with a sequence-to-sequence model [19]. Later on, the focus shifted beyond simple action labels. For example, Guo et al. utilized variational autoencoders to generate motions from text [9] , significantly enhancing the quality and diversity of generated motions. With the success of diffusion models in AIGC, MDM [35] and other related work [42, 16, 5] have introduced diffusion models into the text-to-motion domain, resulting in impressive achievements.
Recently, there has been some work attempting to generate fine-grained or stylized motions. For example, MotionCLIP [34] trains an autoencoder to align motion with the text and images of CLIP in the latent feature space, exploring the generation of stylized actions. MotionDiffuse [42] instead proposes precise control of motions and stylization of body parts using noise interpolation based on the properties of diffusion models. GestureDiffuCLIP [2] incorporates style guidance into the diffusion model through AdaIN [13], enabling the generation of realistic and stylized gestures. For music-to-motion, there has been work [18, 37] trying to generate dance movements of different styles by using different styles of music.
The work most closely related to ours is Action-GPT [15], which introduced, for the first time, Large Language Models (LLMs) into the field of text-conditioned motion generation. Action-GPT can be integrated into any text-to-motion model. However, it enriches only the description of action classes, without providing detailed descriptions of different body parts and further guiding the model training accordingly. For simultaneous action generation, SINC [3] incorporates ChatGPT to identify the body parts involved in the textual description. By generating multiple motions and concatenating them using different body parts, it achieves impressive results. Specifically, SINC divides the human body into [’left arm’, ’right arm’, ’left leg’, ’right leg’, ’torso’, ’neck’, ’buttocks’, ’waist’], which we borrow from in our work. By fine-tuning LLMs, MotionGPT [14, 43] designs a pre-trained motion language model which supports various motion-related tasks through prompts.
2.2 Diffusion Generative Models
The diffusion model is a type of neural generative model based on the stochastic diffusion process in thermodynamics [33, 12]. It starts with samples from the data distribution and gradually adds noise through a forward diffusion process. Then, a neural network learns the reverse process to progressively remove the noise and restore the samples to their original states. Diffusion generative models have achieved significant success in image generation field [30, 31]. For conditional generation,  [6] introduced classifier-guided diffusion, while  [11] proposed a classifier-free method. Given their excellent generation quality,  [35, 42, 16] incorporated diffusion models into motion generation domain, leading to impressive results.
3 Method
Given a textual description, our goal is to generate a human motion $x^{1:n}=\{x^{i}\}_{i=1}^{n}$ that matches the given description. The motion consists of $n$ frames of human poses. For each pose $x^{i}\in\mathbb{R}^{J\times D}$, we represent it by joint rotations or positions, where $J$ represents the number of joints and $D$ represents the dimensionality of the joint representation. In Figure 2, we give an overview of our fine-grained human motion diffusion model. First, we adopt ChatGPT to perform fine-grained paraphrasing of the given vague textual description. This expands concise textual descriptions into descriptions of different body parts. FG-MDM then uses these fine-grained descriptions to guide a diffusion model for human motion generation.
3.1 Prompt Strategy
We first introduce the prompt strategy adopted for generating fine-grained descriptions. For a given textual description of a motion, we utilize ChatGPT-3.5 to generate more fine-grained descriptions based on different body parts. ChatGPT is a conversational model based on a large language model that can engage in natural language conversations and generate corresponding responses. The answers from ChatGPT are often directly influenced by the information and expression provided in the prompt. If the prompt offers clear and detailed questions or instructions, ChatGPT can typically provide relevant and accurate answers. However, if the prompt is too simple, ambiguous, or unclear, ChatGPT may generate unexpected responses or express unclear content. For our task, we carefully design an effective prompt by using experimental verification.
Our designed strategy is: “Translate the motion described by the given sentences to the motion of each body part only using one paragraph. The available body parts include [‘arms’, ‘legs’, ‘torso’, ‘neck’, ‘buttocks’, ‘waist’]. Here are some examples:[Q…A…]. Question: [sentence]”.
3.2 Diffusion Model for Motion Generation
The basic idea of diffusion models [33, 12] is to learn the reverse process of a well-defined stochastic process. Following MDM [35], we design a text-driven human motion generation model based on the diffusion model.
The diffusion model consists of two parts: the forward process and the reverse process, both of which follow the Markov chain. The forward process involves adding noise. The input is the original motion $x_{0}^{1:n}$ from the data distribution, and the output is the motion $x_{t}^{1:n}$ with $t$ times of Gaussian noise added. When enough noise is added, the motion $x_{T}^{1:n}$ can approach the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, aims to reduce the noise in the Gaussian noise $x_{T}^{1:n}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In the denoising process, at diffusion step $t$, a portion of the noise is eliminated, resulting in a less noisy motion $x_{t-1}^{1:n}$. This step is repeated iteratively until the noise is completely removed, generating the denoised motion $\hat{x}_{0}^{1:n}$.
3.2.1 Network
we adopt a simple transformer [38] encoder architecture to implement our network $G$. Unlike the conventional diffusion model mentioned above, we follow [30] and predict the clean motion $\hat{x}_{0}^{1:n}$ instead of predicting the noise added at each diffusion time-step. The input of $G$ is the noised motion $x_{t}^{1:n}$ obtained by adding noise $t$ times to the original motion $x_{0}^{1:n}$. The noised motion $x_{t}^{1:n}$, together with the text condition code $GL$, $PT_{1:k}$ and the noise time-step $t$, is inputted to the transformer encoder, resulting in the clean motion $\hat{x}_{0}^{1:n}$. One of the reasons for directly predicting the clean motion in each step of the diffusion model is to incorporate human geometric losses during the training of the network, making the generated human motions more natural. At each sampling step $t$, from $T$ to 1, our model predicts the clean motion $\hat{x}_{0}^{1:n}$, and then adds noise back to $x_{t-1}^{1:n}$. After $T$ iterations, the final clean motion $\hat{x}_{0}^{1:n}$ is obtained.
3.2.2 Global Token and Part Tokens
For the text condition, we encode the entire fine-grained description with CLIP [29] as a global token $GL$ of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens $PT_{1:k}$, where k denotes the number of body parts. The global token provides overall information about the motion, while part tokens provide local information about different body parts. This design allows the global token and part tokens to jointly guide the diffusion model in generating human motions.
3.2.3 Loss Functions
For training the diffusion model, we follow  [30] to predict the signal itself instead of predicting the noise, i.e., $\hat{x}_{0}^{1:n}\,=\,G(x_{t}^{1:n},t,c)$, with the simple loss function.
$\mathcal{L}_{\mathrm{G}}=E_{x_{0}^{1:n}\sim q(x_{0}^{1:n}|c),t\sim[1,T]}[||x_{0}^{1:n}-G(x_{t}^{1:n},t,c)]|^{2}]$ (1)
In order to generate more natural and kinematically plausible motions, we employ the same geometric losses as MDM [35] from [32, 24], i.e., positions, foot contact, and velocities.
$\mathcal{L}_{\mathrm{pos}}=\frac{1}{n}\sum_{i=1}^{n}||FK(x_{0}^{i})-FK(\hat{x}_{0}^{i})||_{2}^{2},$ (2)
$\mathcal{L}_{\mathrm{foot}}=\frac{1}{n-1}\sum_{i=1}^{n-1}||(FK(\hat{x}_{0}^{i+1})-FK(\hat{x}_{0}^{i}))\cdot f_{i}||_{2}^{2},$ (3)
$\mathcal{L}_{\mathrm{vel}}=\frac{1}{n-1}\sum_{i=1}^{n-1}||(x_{0}^{i+1}-x_{0}^{i})-(\hat{x}_{0}^{i+1}-\hat{x}_{0}^{i})||_{2}^{2}$ (4)
where $FK(\cdot)$ represents the forward kinematic function that converts joint rotations into joint positions. For each frame $i$, $f_{i}\in\{0,1\}^{J}$ is the binary foot contact mask.
Overall, our training loss is
$\mathcal{L}=\mathcal{L}_{\mathrm{G}}+\lambda_{\mathrm{pos}}\mathcal{L}_{\mathrm{pos}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\lambda_{\mathrm{foot}}\mathcal{L}_{\mathrm{foot}}.$ (5)
where $\lambda_{\mathrm{pos}}$, $\lambda_{\mathrm{vel}}$, $\lambda_{\mathrm{foot}}$ are balancing coefficients for the three geometric losses.
4 Experiments
Figure 3: Qualitative results with fine-grained motions. We compare our FG-MDM with MDM [35] and MLD [5]. All three models are trained on HumanML3D. For better visualization, some pose frames are shifted to prevent overlap. Please refer to supplementary materials for more video demos.
Figure 4: Qualitative results with stylized motions. All three models are trained on HumanML3D. Please refer to supplementary materials for more video demos.
[…]
Top 1 Top 2 Top 3
Real $0.520^{\pm.003}$ $0.703^{\pm.006}$ $0.799^{\pm.005}$ $0.004^{\pm.000}$ $2.959^{\pm.007}$ $9.412^{\pm.046}$
T2M-GPT [41] $0.269^{\pm.008}$ $0.410^{\pm.009}$ $0.507^{\pm.012}$ $3.910^{\pm.069}$ $5.014^{\pm.060}$ $8.128^{\pm.171}$ $2.578^{\pm.122}$
MLD [5] $0.286^{\pm.002}$ $0.443^{\pm.003}$ $0.552^{\pm.003}$ $2.244^{\pm.053}$ $4.725^{\pm.015}$ $8.773^{\pm.058}$ $\textbf{3.317}^{\pm.078}$
MotionDiffuse [42] $0.308^{\pm.007}$ $0.466^{\pm.008}$ $0.571^{\pm.007}$ $2.755^{\pm.087}$ $\textbf{4.483}^{\pm.031}$ $8.333^{\pm.163}$ $1.667^{\pm.127}$
MDM [35] $0.187^{\pm.009}$ $0.309^{\pm.007}$ $0.399^{\pm.008}$ $4.363^{\pm.265}$ $7.331^{\pm.065}$ $8.432^{\pm.117}$ $3.059^{\pm.027}$
FG-MDM $\textbf{0.308}^{\pm.015}$ $\textbf{0.484}^{\pm.005}$ $\textbf{0.593}^{\pm.003}$ $\textbf{0.663}^{\pm.051}$ $5.649^{\pm.035}$ $\textbf{9.476}^{\pm.077}$ $2.632^{\pm.085}$
Table 1: Quantitative results on HumanML3D. To examine the ability in handling fine-grained descriptions, for all the four control methods, we employ their pre-trained models to generate motions by using our fine-grained description as input.
[…]
Top 1 Top 2 Top 3
Real $0.418^{\pm.011}$ $0.633^{\pm.011}$ $0.775^{\pm.009}$ $0.038^{\pm.009}$ $2.749^{\pm.034}$ $10.852^{\pm.194}$
MotionDiffuse [42] $\textbf{0.189}^{\pm.008}$ $\textbf{0.296}^{\pm.016}$ $\textbf{0.376}^{\pm.009}$ $16.136^{\pm.648}$ $\textbf{6.513}^{\pm.104}$ $8.948^{\pm.339}$ $1.024^{\pm.027}$
MDM [35] $0.099^{\pm.013}$ $0.167^{\pm.013}$ $0.235^{\pm.019}$ $16.372^{\pm.855}$ $10.501^{\pm.058}$ $8.758^{\pm.233}$ $\textbf{2.242}^{\pm.998}$
FG-MDM $0.151^{\pm.010}$ $0.269^{\pm.013}$ $0.363^{\pm.008}$ $\textbf{0.344}^{\pm.016}$ $9.351^{\pm.073}$ $\textbf{10.707}^{\pm.082}$ $0.871^{\pm.442}$
Table 2: Quantitative results on KIT. Since T2M-GPT [41] and MLD [5] do not release their pre-trained models on KIT. We thus compare our model with MotionDiffuse and MDM only.
In this section, we first elaborate the datasets, evaluation metrics, and implementation details. We then conduct quantitative experiments to compare FG-MDM with current state-of-the-art approaches. In order to completely evaluate FG-MDM, we further present two ablation studies. To show the generalization capability of our model, we further perform qualitative experiments and a user study to examine FG-MDM’s ability in generating fine-grained and stylized motions.
4.1 Experimental Details
4.1.1 Datasets
We utilize the HumanML3D [9] dataset and the KIT [26] dataset to train and evaluate our model. HumanML3D is a recently proposed large-scale dataset of motion-text pairs. It consists of 14,616 motion sequences from the AMASS [21] and HumanAct12 [8] datasets, with multiple ways of describing each motion, resulting in a total of 44,970 text annotations. The KIT dataset, on the other hand, is relatively smaller and contains 3,911 motion sequences along with their corresponding 6,353 text descriptions. For both datasets, we use 80% of the data for training, and use the remaining for test.
We preprocess the 44,970 text descriptions from HumanML3D and 6,353 text descriptions from KIT using ChatGPT-3.5. This extends these descriptions into fine-grained ones for our model training.
4.1.2 Evaluation Metrics
For quantitative experiments, we employ five evaluation metrics to evaluate our model’s ability in fitting the training data: R Precision, FID, Multimodal Dist, Diversity, and Multimodality. R Precision and Multimodal Dist are used to assess the correlation between generated motions and input text. Diversity and Multimodality are utilized to evaluate the diversity of generated motions. FID measures the difference in feature distribution between generated motions and ground truth in latent space, which is utilized to evaluate the quality of generated motions.
4.1.3 Implementation Details
In our study, the transformer accepts tokens whose feature dimension is 512 as input. We use 4 attention heads and apply a dropout rate of 0.1. The transformer encoder consists of 8 stacked encoder layers to capture complex relationships and hierarchies in the data. For ChatGPT, we adopt the gpt-3.5-turbo API provided by OpenAI. For text encoding, we employ the frozen CLIP-ViT-B/32 model as the encoder. Our batch size is set to 64. Additionally, we set the diffusion step to 1000. On a single NVIDIA GeForce RTX3090 GPU, it takes about 6 days for training our model.
4.2 Comparison with Prior Work
To evaluate the performance of FG-MDM in handling fine-grained description-based motion generation, we compare our work with four recent motion generation approaches: T2M-GPT [41], MLD [5], MotionDiffuse [42], and MDM [35]. Note that all these four methods achieve state-of-the-art performance when using short text as input. To examine their ability in handling fine-grained descriptions, for all these four methods, we employ their pre-trained models to generate motions by using our fine-grained description as input. In Table 1 and Table 2, we provide experimental results on the HumanML3D and KIT datasets, respectively. Note that on KIT, T2M-GPT and MLD do not release their pre-trained models. We thus compare our model with MotionDiffuse and MDM only on KIT. For all experiments, We run the evaluation 5 times and “±” indicates the 95% confidence interval.
As reported in both tables, we achieve impressive performance when handling the task of fine-grained description-based motion generation. In particular, FG-MDM leads to much shorter FID than prior state-of-the-art. For example, on KIT, both MotionDiffuse and MDM get long FID around 16, while our model gets a figure shorter than 0.35. This indicates FG-MDM can generate motions closer to real motions than other methods, when using fine-grained description as input.
4.3 Ablation Study
In order to validate the contribution of different modules, next we conduct two ablation studies. The first study is used to examine the contribution of ChatGPT-Generated fine-grained texts, which is performed by replacing the fine-grained description with the original short text. Therefore, only a single token is adopted to guide the diffusion model. The second study is performed by adopting only the global token to check the contribution of part tokens when fine-grained descriptions are used. In Table 3 and Table 4, we give the results on HumanML3D and KIT, respectively. As observed, ablating either fine-grained descriptions or part tokens leads to poorer performance.
4.4 Generalization Capability Study
Note that FG-MDM aims to not only fit the training data but also generate fine-grained and stylized motions beyond the scope of the dataset. Considering that HumanML3D and KIT include very few fine-grained and stylized motions, we customize some texts that are not in the two datasets to show the effectiveness of FG-MDM in generating fine-grained and stylized motions. In Figure 3 and Figure 4, we show motions generated by MDM [35], MLD [5] and our FG-MDM. Note that for all the three methods, we use models trained on HumanML3D to generate motions. In comparison, our method generates motions that are more consistent with the details described in the slightly complex textual descriptions. Even for difficult stylized textual descriptions, our method is capable of generating motions that correspond to the respective styles. This demonstrates the strong generalization capability of FG-MDM. Please refer to supplementary materials for more video demos.
Figure 5: User study results. For each method, a color bar ranging from blue to red is used to represent the percentage of text-to-motion match levels, with blue indicating the least match and red indicating the most match.
4.5 User Study
To further examine FG-MDM’s generalization capability, we conduct a user study to evaluate the quality of motions generated by our model based on human visual perception. We customize a total of 40 fine-grained and stylized textual descriptions. With these descriptions, we generate motions by using MDM [35], MotionDiffuse [42] and our FG-MDM. We then recruit 10 users for the study. In each question, participants are asked to rate the degree of matching between the generated motion and the textual description on a scale of 0 to 2. The results are given in Figure 5. Overall, users believe that FG-MDM matches texts much better in generating fine-grained and stylized motions than the other two methods. For example, the proportion of MDM and MotionDiffuse obtaining the most matching scores is only slightly higher than half of FG-MDM.
5 Conclusion
In this study, we use ChatGPT to perform fine-grained paraphrasing on the textual annotations of HumanML3D and KIT. With these fine-grained descriptions, We explore a Fine-Grained Human Motion Diffusion Model. It utilizes fine-grained descriptions of different body parts to guide the training of a diffusion model. This enables it to learn the generation of fine-grained and stylized motions even outside of the distribution of the training dataset. In the future, we would like to improve the quality of fine-grained annotations of human motions.
[…]
$4.363$ $0.399$ $8.432$
$1.050$ $0.427$ $9.509$
0.663 0.593 9.476
Table 3: Ablation study results on HumanML3D. “F” denotes using ChatGPT-generated fine-grained descriptions. “P” represents adopting part tokens.
[…]
$16.372$ $0.235$ $8.758$
$0.549$ $0.360$ 10.829
0.344 0.383 $10.707$
Table 4: Ablation study results on KIT.